Given {2.4, 3.9, 5.4, 6.9, 8.4, ...}

(a) Find the 111th term.

(b) Find the sum of the first 111 terms of the sequence.



1
Expert's answer
2021-02-03T03:44:08-0500

"a=2.4"


"d=3.9-2.4=5.4-3.9=6.9-5.4"

"=8.4-6.9=1.5"

a)

"a_{111}=a+110d=2.4+110(1.5)=167.4"



b)


"S_{111}=111\cdot\dfrac{a+a_{111}}{2}=111\cdot\dfrac{2.4+167.4}{2}"

"=9423.9"